# Structure-preserving contrastive learning for spatial time series

### v3

@@ Abstract
Neural network models are increasingly applied in transportation research to tasks such as prediction. The effectiveness of these models largely relies on learning meaningful latent patterns from data, where self-supervised learning of informative representations can enhance model performance and generalisability. However, self-supervised representation learning for spatially characterised time series, which are ubiquitous in transportation domain, poses unique challenges due to the necessity of maintaining fine-grained spatio-temporal similarities in the latent space. In this study, we introduce two structure-preserving regularisers for the contrastive learning of spatial time series: one regulariser preserves the topology of similarities between instances, and the other preserves the graph geometry of similarities across spatial and temporal dimensions. To balance the contrastive learning objective and the need for structure preservation, we propose a dynamic weighting mechanism that adaptively manages this trade-off and stabilises training. We validate the proposed method through extensive experiments, including multivariate time series classification to demonstrate its general applicability, as well as macroscopic and microscopic traffic prediction to highlight its particular usefulness in encoding traffic interactions. Across all tasks, our method preserves the similarity structures more effectively and improves state-of-the-art task performances. This method can be integrated with an arbitrary neural network model and is particularly beneficial for time series data with spatial or geographical features. Furthermore, our findings suggest that well-preserved similarity structures in the latent space indicate more informative and useful representations. This provides insights to design more effective neural networks for data-driven transportation research. Our code is made openly accessible with all resulting data at https://github.com/yiru-jiao/spclt.
@@ 1 Introduction
The proposed method is applicable to spatial time series in general, while we highlight its particular usage for traffic interactions in this paper. To thoroughly validate the method, we conduct experiments on tasks of 1) multivariate time series classification, where we benchmark against the current state-of-the-art (SOTA) models, i.e., [16] and [17]; and 2) traffic prediction, where we use [18] for macroscopic benchmark and [19] for microscopic. Along with these experiments, the efficiency of this method is evaluated with various network architectures. In addition to performance improvement, we also investigate and discuss the importance of preserving similarity structure. Below is a summary of the key contributions in this study.
@@
  • Considering neural network modelling as learning the conditional probability distribution of outputs over inputs, the similarity structure hidden in the input data implies the distribution of conditions. Our method is therefore important to preserve the original distribution in the latent space for more effective representational learning. This is particularly beneficial for time series data with spatial or geographical features as they require finer-grained information, not only in the field of transportation, but also robotics, meteorology, remote sensing, etc.
@@ 2.1 Time series contrastive learning
Contrastive learning for time series data is a relatively young niche and is rapidly developing. The development has been dominantly focused on defining positive and negative samples. Early approaches construct positive and negative samples with subseries within time series [20] and temporal neighbourhoods [21]; and later methods create augmentations by transforming original series [22, 23]. More recently, [16] generates random masks to enable both instance-wise and time-wise contextual representations at flexible hierarchical levels, which exceeds previous state-of-the-art performances (SOTAs). Given that not all negatives may be useful [24, 25], [26] makes hard negatives to boost performance, while [17] utilises soft contrastive learning to weigh sample pairs of varying similarities, both of which reach new SOTAs.
@@ 2.2 Structure-preserving SSRL
Preserving the original structure of data when mapping into a latent space has been widely and actively researched in manifold learning (for a literature review, see [28]) and graph representation learning [29, 30]. In manifold learning, which is also known as nonlinear dimension reduction, the focus is on revealing the geometric shape of data point clouds for visualisation, denoising, and interpretation. In graph representation learning, the focus is on maintaining the connectivity of nodes in the graph while compressing the data space required for large-scale graphs [31]. Structure-preserving has not yet attracted much dedication to time series data. [32] provides a literature review on time series data dimensionality reduction, where none of the methods are specifically tailored for time series.
@@ 2.3 Traffic interaction SSRL
In line with the conclusions in previous sub-sections, existing exploration in the context of traffic interaction data and tasks also predominantly relies on autoencoders and graphs. For instance, using a transformer-based multivariate time series autoencoder [41], [42] clusters traffic scenarios with trajectories of pairwise vehicles. Then a series of studies investigate masking strategies with autoencoders for individual trajectories and road networks, including [43, 44, 45].
Leveraging data augmentation, [46] utilises graphs and contrastive learning to jointly learn representations for vehicle trajectories and road networks. The authors design road segment positive samples as neighbours in the graph, and trajectory positive samples by replacing a random part with another path having the same origin and destination. Similarly, [47] uses a graph-based contrastive learning approach to learn traffic scene similarity. The authors randomly modify the position and velocity of individual traffic participants in a scene to construct positive samples, with negative samples drawn uniformly from the rest of a training batch. Also using augmentation, [48] focuses on capturing seasonal and holiday information for traffic prediction.
@@ 3 Methods
This section begins by defining the problem of structure-preserving contrastive learning for spatial time series. Following that, we explain the overall loss function to be optimised, where we propose a dynamic weighing mechanism to balance contrastive learning and structure preservation during training. Then we present the contrastive learning loss for time series, unifying both hard and soft versions in a consistent format. Lastly, we introduce two structure-preserving regularisers, which are respectively adapted to maintain the global and local structure of similarity relations.
@@ 3.1 Problem definition
3.2 Trade-off between contrastive learning and structure preservation
We define the complete loss function for optimising $f_{\bm{\theta}}$ as shown in Equation (1). Referring to the simplified loss in Section 1, i.e., ${\mathcal{L}}=\eta_{\text{CLT}}\cdot\ell_{\text{CLT}}+\eta_{\text{SP}}\cdot\ell_{\text{SP}}+r_{\bm{\eta}}$, the contrastive learning loss for time series (${\mathcal{L}}_{\text{CLT}}$) and structure-preserving loss (${\mathcal{L}}_{\text{SP}}$) are modified using the function $x(1-\exp(-x))$ and correspond to $\ell_{\text{CLT}}$ and $\ell_{\text{SP}}$; $\eta_{\text{CLT}}$, $\eta_{\text{SP}}$, and $r_{\bm{\eta}}$ depend on two deviation terms $\sigma_{\text{CLT}}$ and $\sigma_{\text{SP}}$, which dynamically change during training.
@@ 3.2 Trade-off between contrastive learning and structure preservation
The modification by $x(1-\exp(-x))$ serves two purposes: it penalises negative values of ${\mathcal{L}}_{\text{SP}}$, and stabilises training when either ${\mathcal{L}}_{\text{CLT}}$ or ${\mathcal{L}}_{\text{SP}}$ approaches its optimal value. While in computation the value of ${\mathcal{L}}_{\text{SP}}$ sometimes is below zero, the losses used as ${\mathcal{L}}_{\text{CLT}}$ and ${\mathcal{L}}_{\text{SP}}$ in this study all have their theoretical optimal values of zero [^1] We offer a more detailed analysis in A.. As ${\mathcal{L}}_{\text{CLT}}$ decreases and approaches its optimum zero, the modified term $\ell_{\text{CLT}}$ has a slower decreasing rate when ${\mathcal{L}}_{\text{CLT}}<1$. More specifically, the derivative of $x(1-\exp(-x))$ is $x^{\prime}(1-\exp(-x)(1-x))$, where $x^{\prime}$ denotes the derivative of $x$ and the multiplier in parentheses decreases from 1 to 0 while $x$ decreases from 1 to 0. This thus stabilises the training when ${\mathcal{L}}_{\text{CLT}}$ approaches zero, and works the same for ${\mathcal{L}}_{\text{SP}}$.
Inspired by [49], we then weigh the two modified losses by considering their uncertainties. The magnitudes of ${\mathcal{L}}_{\text{CLT}}$ and ${\mathcal{L}}_{\text{SP}}$ may vary with different datasets and hyperparameter settings. This variation precludes fixed weights for contrastive learning and structure preservation. We consider the loss values (denoted by $\ell$) as deviations from their optimal values, and learn adaptive weights according to the deviations. Given the optimal value of 0, we assume a Gaussian distribution of $\ell$ with standard deviation $\sigma$, i.e., $p(\ell)={\mathcal{N}}(0,\sigma^{2})$. Then we can maximise the log likelihood $\sum\log p(\ell)=\sum(-\log 2\pi-\log\sigma^{2}-\ell^{2}/\sigma^{2})/2$ to learn $\sigma$. This is equivalent to minimising $\sum\left(\ell^{2}/2/\sigma^{2}+\log\sigma\right)$. When balancing between two losses $\ell_{\text{CLT}}$ and $\ell_{\text{SP}}$ that have deviations $\sigma_{\text{CLT}}$ and $\sigma_{\text{SP}}$, respectively, we need to use Equation (2).
Replacing $\ell_{\text{CLT}}$ in Equation (2) with ${\mathcal{L}}_{\text{CLT}}\left(1-\exp(-{\mathcal{L}}_{\text{CLT}})\right)$ and $\ell_{\text{SP}}$ with ${\mathcal{L}}_{\text{SP}}\left(1-\exp(-{\mathcal{L}}_{\text{SP}})\right)$, Equation (1) is then derived to be the complete loss. The training process trades-off between ${\mathcal{L}}_{\text{CLT}}$ and ${\mathcal{L}}_{\text{SP}}$, as well as between the weight regulariser $r_{\bm{\eta}}=\log\sigma_{\text{CLT}}\sigma_{\text{SP}}$ and the rest of Equation (1). When ${\mathcal{L}}_{\text{CLT}}$ is small and ${\mathcal{L}}_{\text{SP}}$ is large, $\sigma_{\text{CLT}}$ becomes small and $\sigma_{\text{SP}}$ becomes large, which then increases the weight for ${\mathcal{L}}_{\text{CLT}}$ while reduces the weight for ${\mathcal{L}}_{\text{SP}}$. The reverse occurs when ${\mathcal{L}}_{\text{CLT}}$ is large and ${\mathcal{L}}_{\text{SP}}$ is small. As the weighted sum of ${\mathcal{L}}_{\text{CLT}}$ and ${\mathcal{L}}_{\text{SP}}$ increases by larger weights, $\log\sigma_{\text{CLT}}\sigma_{\text{SP}}$ decreases and discourages the increase from being too much. Similarly, if the weighted sum decreases by smaller weights, $\log\sigma_{\text{CLT}}\sigma_{\text{SP}}$ also regularises the decrease.
3.3 Contrastive learning loss
@@
In this study, we use the time series contrastive learning loss introduced in TS2Vec [16] and its succeeder SoftCLT [17] that utilises soft weights for similarity comparison [^2] We write the loss function equations in this sub-section following the code open-sourced with the original papers; as such, they are slightly adjusted from those in the original papers.. For each sample ${\bm{x}}_{i}$, two augmentations are created by timestamp masking and random cropping, and then encoded as two representations ${\bm{z}}_{i}^{\prime}$ and ${\bm{z}}_{i}^{\operatorname*{{\prime\prime}}}$. TS2Vec and SoftCLT losses consider the same sum of similarities for a sample $i$ at a timestamp $t$, as shown in Equations (3) and (4). Equation (3) is used for instance-wise contrasting, which we denote by the subscript ${}_{\text{inst}}$; Equation (4) is used for time-wise contrasting, denoted by the subscript ${}_{\text{temp}}$.
@@
where $\begin{cases}\displaystyle\ell_{\begin{subarray}{c}\text{inst}\\ \text{TS2Vec}\end{subarray}}^{(i,t)}=-\log\frac{\exp({\bm{z}}^{\prime}_{i,t}\cdot{\bm{z}}^{\operatorname*{{\prime\prime}}}_{i,t})+\exp({\bm{z}}^{\operatorname*{{\prime\prime}}}_{i,t}\cdot{\bm{z}}^{\prime}_{i,t})}{S_{\text{inst}}(i,t)}\\ \displaystyle\ell_{\begin{subarray}{c}\text{temp}\\ \text{TS2Vec}\end{subarray}}^{(i,t)}=-\log\frac{\exp({\bm{z}}^{\prime}_{i,t}\cdot{\bm{z}}^{\operatorname*{{\prime\prime}}}_{i,t})+\exp({\bm{z}}^{\operatorname*{{\prime\prime}}}_{i,t}\cdot{\bm{z}}^{\prime}_{i,t})}{S_{\text{temp}}(i,t)}\end{cases}$
Similarity comparison in TS2Vec is between two different augmentations for the same sample. This is expanded by SoftCLT to also involve other samples in the same mini-batch. Varying instance-wise and time-wise weights are assigned to different comparison pairs as soft assignments, with Equations (6) and (7). This introduces four hyperparameters, i.e., $\tau_{\text{inst}}$, $\tau_{\text{temp}}$, $\alpha$, and $m$. We use DTW to compute $d({\bm{x}}_{i},{\bm{x}}_{j})$ and set $\alpha=0.5$, as recommended in the original paper; the other parameters need to be tuned for different datasets. Specifically, $m$ controls the sharpness of time hierarchical contrasting. TS2Vec uses $m=1$ (constant) and SoftCLT uses $m(k)=2^{k}$ (exponential), where $k$ is the depth of pooling layers when computing temporal loss. In this study, we add one more option $m(k)=k+1$ (linear), and will tune the best way for different datasets.
@@
where $\ell_{\begin{subarray}{c}\text{inst}\\ \text{SoftCLT}\end{subarray}}^{(i,t)}&=-\sum_{j=1}^{B}w_{\text{inst}}(i,j)\log\frac{\exp({\bm{z}}^{\prime}_{i,t}\cdot{\bm{z}}^{\operatorname*{{\prime\prime}}}_{j,t})+\exp({\bm{z}}^{\operatorname*{{\prime\prime}}}_{i,t}\cdot{\bm{z}}^{\prime}_{j,t})}{S_{\text{inst}}(i,t)}\\ &\quad-\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{B}w_{\text{inst}}(i,j)\log\frac{\exp({\bm{z}}^{\prime}_{i,t}\cdot{\bm{z}}^{\prime}_{j,t})+\exp({\bm{z}}^{\operatorname*{{\prime\prime}}}_{i,t}\cdot{\bm{z}}^{\operatorname*{{\prime\prime}}}_{j,t})}{S_{\text{inst}}(i,t)}\end{aligned}\\ \begin{aligned} \ell_{\begin{subarray}{c}\text{temp}\\ \text{SoftCLT}\end{subarray}}^{(i,t)}&=-\sum_{s=1}^{T}w_{\text{temp}}(t,s)\log\frac{\exp({\bm{z}}^{\prime}_{i,t}\cdot{\bm{z}}^{\operatorname*{{\prime\prime}}}_{i,s})+\exp({\bm{z}}^{\operatorname*{{\prime\prime}}}_{i,t}\cdot{\bm{z}}^{\prime}_{i,s})}{S_{\text{temp}}(i,t)}\\ &\quad-\sum_{\begin{subarray}{c}s=1\\ s\neq t\end{subarray}}^{T}w_{\text{temp}}(t,s)\log\frac{\exp({\bm{z}}^{\prime}_{i,t}\cdot{\bm{z}}^{\prime}_{i,s})+\exp({\bm{z}}^{\operatorname*{{\prime\prime}}}_{i,t}\cdot{\bm{z}}^{\operatorname*{{\prime\prime}}}_{i,s})}{S_{\text{temp}}(i,t)}\end{aligned}\end{cases}$
@@ 3.4 Structure-preserving regularisers
The graph-geometry-preserving loss is also computed per mini-batch, as is shown in Equation (10). This loss measures geometry distortion, i.e., how much $f_{\bm{\theta}}$ deviates from being an isometry that preserves distances and angles. The geometry to be preserved of the original data manifold is implied by a similarity graph. To represent temporal and spatial characteristics, instead of using an instance as a node in the graph, we consider the nodes as timestamps or in a spatial dimension such as sensors or road users. Then the edges in the graph are defined by pairwise geodesic distances between nodes.

### v4

@@ Abstract
Neural network models are increasingly applied in transportation research to tasks such as prediction. The effectiveness of these models largely relies on learning meaningful latent patterns from data, where self-supervised learning of informative representations can enhance model performance and generalisability. However, self-supervised representation learning for spatially characterised time series, which are ubiquitous in transportation domain, poses unique challenges due to the necessity of maintaining fine-grained spatio-temporal similarities in the latent space. In this study, we introduce two structure-preserving regularisers for the contrastive learning of spatial time series: one regulariser preserves the topology of similarities between instances, and the other preserves the graph geometry of similarities across spatial and temporal dimensions. To balance the contrastive learning objective and the need for structure preservation, we propose a dynamic weighting mechanism that adaptively manages this trade-off and stabilises training. We validate the proposed method through extensive experiments, including multivariate time series classification to demonstrate its general applicability, as well as macroscopic and microscopic traffic prediction to highlight its particular usefulness in encoding traffic interactions. Across all tasks, our method preserves the similarity structures more effectively and improves state-of-the-art task performances. This method can be integrated with an arbitrary neural network model and is particularly beneficial for time series data with spatial or geographical features. Furthermore, our findings suggest that well-preserved similarity structures in the latent space indicate more informative and useful representations. This provides insights to design and optimise more effective neural networks for data-driven transportation research. Our code is made openly accessible with all resulting data at https://github.com/yiru-jiao/spclt.
@@ 1 Introduction
The proposed method is applicable to spatial time series in general, while we highlight its particular usage for traffic interactions in this paper. To thoroughly validate the method, we conduct experiments on tasks of 1) multivariate time series classification, where we benchmark against the current state-of-the-art (SOTA) models, i.e., [16] and [17]; and 2) traffic prediction, where we use [18] for macroscopic benchmark and [19] for microscopic. Along these experiments, the efficiency of this method is evaluated with multiple network architectures. In addition to performance improvement, we also investigate the impacts of preserving similarity structure during training. Below is a summary of the contributions in this study.
@@
  • Considering neural network modelling as learning the conditional probability distribution of outputs over inputs, the similarity structure hidden in the input data implies the distribution of conditions. Our method is therefore important to preserve the original distribution in the latent space for more effective model training. This is particularly beneficial when dealing with spatial time series data in transportation domain, where fine-grained and hierarchical information is required in modelling.
@@ 2.1 Time series contrastive learning
Contrastive learning for time series data is a relatively young niche and is rapidly developing. The development has been dominantly focused on defining positive and negative samples. Early approaches construct positive and negative samples with subseries within time series [20] and temporal neighbourhoods [21]; and later methods create augmentations by transforming original series [22, 23]. More recently, [16] generates random masks to enable both instance-wise and time-wise contextual representations at flexible hierarchical levels, which exceeds previous state-of-the-art performances (SOTAs). Given that not all negatives may be useful [24, 25][26] makes hard negatives to boost performance, while [17] utilises soft contrastive learning to weigh sample pairs of varying similarities, both of which reach new SOTAs.
@@ 2.2 Structure-preserving SSRL
Preserving the original structure of data when mapping into a latent space has been widely and actively researched in manifold learning (for a literature review, see [28]) and graph representation learning [29, 30]. In manifold learning, which is also known as nonlinear dimension reduction, the focus is on revealing the geometric shape of data point clouds for visualisation, denoising, and interpretation. In graph representation learning, the focus is on maintaining the connectivity of nodes in the graph while compressing the data space required for large-scale graphs [31]. Structure-preserving has not yet attracted much dedication to time series data. [32] provides a literature review on time series data dimensionality reduction, where none of the methods are specifically tailored for time series. Existing studies that are the most relevant include [33, 34, 35], which construct hierarchies of samples or features, while similarity preservation remains under-explored.
@@ 2.3 Traffic interaction SSRL
In line with the literature summary in previous sub-sections, existing exploration of SSRL in the context of traffic interaction data and tasks have been predominantly relied on autoencoders and graph-based contrastive learning. For instance, using a transformer-based multivariate time series autoencoder [44],  [45] clusters traffic scenarios with trajectories of pairwise vehicles. Then a series of studies investigate masking strategies with autoencoders for individual trajectories and road networks, including [46, 47, 48]. Combining graph (convolutional) neural networks and contrastive learning, a variety of studies have shown accuracy and stability improvements in traffic flow prediction [49, 50, 51, 52].
There are some other use cases. Leveraging data augmentation, [53] utilises graphs and contrastive learning to jointly learn representations for vehicle trajectories and road networks. The authors design road segment positive samples as neighbours in the graph, and trajectory positive samples by replacing a random part with another path having the same origin and destination. In a similar way, [54] learns traffic scene similarity. The authors randomly modify the position and velocity of individual traffic participants in a scene to construct positive samples, with negative samples drawn uniformly from the rest of a training batch. By designing augmentation based on domain-knowledge, [55] focuses on capturing seasonal and holiday information for traffic prediction, while [56] targets abnormal traffic patterns caused by incidents or reduced capacities.
@@ 3.1 Problem definition
3.2 SPCLT loss
Equation (1) presents an overview of the structure-preserving contrastive learning loss for time series, abbreviated as SPCLT loss, to optimise $f_{\bm{\theta}}$ for self-supervised representation learning.
@@ 3.2 SPCLT loss
Referring to the simplified loss in Section 1, i.e., ${\mathcal{L}}=\eta_{\text{CLT}}\cdot\ell_{\text{CLT}}+\eta_{\text{SP}}\cdot\ell_{\text{SP}}+r_{\bm{\eta}}$, the contrastive learning loss for time series (${\mathcal{L}}_{\text{CLT}}$) and structure-preserving loss (${\mathcal{L}}_{\text{SP}}$) are modified using the function $y=x(1-\exp(-x))$ and correspond to $\ell_{\text{CLT}}$ and $\ell_{\text{SP}}$. This modification is designed to stabilise ${\mathcal{L}}_{\text{CLT}}$ and ${\mathcal{L}}_{\text{SP}}$ when they are close to 0. The terms $\eta_{\text{CLT}}$, $\eta_{\text{SP}}$, and $r_{\bm{\eta}}$ controls the trade-off between contrastive leraning and structure preservation, depending on two deviations $\sigma_{\text{CLT}}$ and $\sigma_{\text{SP}}$ that dynamically change during training.
In the following sub-sections, we will first introduce the component losses for time series contrastive learning and structure preservation, and then provide more detailed explanations on their stabilisation and dynamic trade-offs.
3.3 Contrastive learning for time series
@@
In this study, we use the time series contrastive learning loss introduced in TS2Vec [16] and its succeeder SoftCLT [17] that utilises soft weights for similarity comparison [^1] We unify the loss function equations in a consistent format following the open-source code provided with the original papers; as such, they are slightly adjusted from the equations in the original papers.. For each sample ${\bm{x}}_{i}$, two augmentations are created by timestamp masking and random cropping, and then encoded as two representations ${\bm{z}}_{i}^{\prime}$ and ${\bm{z}}_{i}^{\operatorname*{{\prime\prime}}}$. TS2Vec and SoftCLT losses consider the same sum of similarities for a sample $i$ at a timestamp $t$, as shown in Equations (2) and (3). Equation (2) is used for instance-wise contrasting, which we denote by the subscript $_{\text{inst}}$; Equation (3) is used for time-wise contrasting, denoted by the subscript $_{\text{temp}}$.
@@
[…]
Similarity comparison in TS2Vec is between two different augmentations for the same sample. This is expanded by SoftCLT to also involve other samples in the same mini-batch. Varying instance-wise and time-wise weights are assigned to different comparison pairs as soft assignments, with Equations (5) and (6). This introduces four hyperparameters, i.e., $\tau_{\text{inst}}$, $\tau_{\text{temp}}$, $\alpha$, and $m$. We use EUC to compute $d({\bm{x}}_{i},{\bm{x}}_{j})$ throughout this paper for convenience, and set $\alpha=0.5$ as recommended in the original paper; the other parameters need to be tuned for different datasets. Specifically, $m$ controls the sharpness of time hierarchical contrasting. TS2Vec uses $m=1$ (constant) and SoftCLT uses $m(k)=2^{k}$ (exponential), where $k$ is the depth of pooling layers when computing temporal loss. In this study, we add one more option $m(k)=k+1$ (linear), and will tune the best way for different datasets.
@@
[…]
@@ 3.4 Structure-preserving regularisers
The graph-geometry-preserving loss is also computed per mini-batch, as is shown in Equation (9). Note that the loss in [41] omits $P$, while we retain it so that ${\mathcal{L}}_{\text{GGeo}}$ measures geometry distortion, i.e., how much $f_{\bm{\theta}}$ deviates from being an isometry that preserves distances and angles. The geometry to be preserved of the original data manifold is implied by a similarity graph. To represent temporal and spatial characteristics, instead of using an instance as a node in the graph, we consider the nodes as timestamps or in a spatial dimension such as sensors or road users. Then the edges in the graph are defined by pairwise geodesic distances between nodes.
@@
3.5 Stabilisation around the theoretical optimal values
This study considers ${\mathcal{L}}_{\text{TS2Vec}}$ and ${\mathcal{L}}_{\text{SoftCLT}}$ as ${\mathcal{L}}_{\text{CLT}}$, and ${\mathcal{L}}_{\text{SP}}$ can be ${\mathcal{L}}_{\text{Topo}}$ or ${\mathcal{L}}_{\text{GGeo}}$. Under this consideration, the optimal values for both ${\mathcal{L}}_{\text{CLT}}$ and ${\mathcal{L}}_{\text{SP}}$ are 0. For ${\mathcal{L}}_{\text{TS2Vec}}$, a value of 0 is reached when ${\bm{z}}_{i,t}^{\prime}$ and ${\bm{z}}_{i,t}^{\operatorname*{{\prime\prime}}}$ are identical. Similarly, the optimal case of ${\mathcal{L}}_{\text{SoftCLT}}$ is when the samples with soft assignments close to 1 are identical, while dissimilar samples have soft assignments close to 0. The topology-preserving loss ${\mathcal{L}}_{\text{Topo}}$ is 0 when the topologically relevant distances remain the same in the latent space as in the original space, i.e., ${\bm{A}}^{X}\left[\bm{\pi}^{X}\right]={\bm{A}}^{Z}\left[\bm{\pi}^{X}\right]$ and ${\bm{A}}^{X}\left[\bm{\pi}^{Z}\right]={\bm{A}}^{Z}\left[\bm{\pi}^{Z}\right]$. Finally, ${\mathcal{L}}_{\text{GGeo}}$ approximates the distortion measure of isometry and is ideally 0, although it may be negative as the approximation of $\tilde{H}_{i}$ is kernel-based depending on the hyperparameter $h$.
Figure 1: Illustration of the modification function.
Section 3.2 has briefly described the modification of ${\mathcal{L}}_{\text{CLT}}$ and ${\mathcal{L}}_{\text{SP}}$ by $y=x(1-\exp(-x))$ in Equation (1). Now we explain the effect of this function more with Figure 1. The first objective of this modification is to penalise negative values of ${\mathcal{L}}_{\text{CLT}}$ and ${\mathcal{L}}_{\text{SP}}$. While $x>0$, $y$ decreases as $x$ decreases; but when $x<0$, $y$ rapidly increases as $x$ decreases. Once ${\mathcal{L}}<0$, the modified term $\ell$ increases and the direction of gradient descent reverses.
Another objective is to avoid large-step updates when $\mathcal{L}$} is close to 0, so as not to miss its optimum. The modified $y$ approximates $x$ while $x$ is large, but has a slower decreasing rate when $x<1$. More specifically, the derivative of $x(1-\exp(-x))$ is $x^{\prime}(1-\exp(-x)(1-x))$, where $x^{\prime}$ denotes the derivative of $x$. While $x>1$, the multiplier in the parentheses is around 1 and $x^{\prime}$ is less interfered with. While $x$ decreases from 1 to 0, the multiplier decreases from 1 to 0, so the derivative also decreases. As a result, this modification can stabilise training when either ${\mathcal{L}}_{\text{CLT}}$ or ${\mathcal{L}}_{\text{SP}}$ approaches its optimal value 0.
3.6 Dynamic weighting mechanism to balance contrastive learning and structure preservation
The training needs to balance between contrastive learning and structure preservation to avoid the neural network parameters being biased by either of the two objectives. However, the magnitudes of ${\mathcal{L}}_{\text{CLT}}$ and ${\mathcal{L}}_{\text{SP}}$ vary with different datasets and hyperparameter settings. This variation precludes fixed weights for the two modified losses.
Inspired by [57], we weigh $\ell_{\text{CLT}}$ and $\ell_{\text{SP}}$ by considering their uncertainties. We consider the loss values as deviations from their optimal values, and learn adaptive weights according to the deviations. Given the optimal value of 0, we assume a Gaussian distribution of $\ell$ with standard deviation $\sigma$, i.e., $p(\ell)={\mathcal{N}}(0,\sigma^{2})$. Then we can maximise the log likelihood $\sum\log p(\ell)=\sum(-\log 2\pi-\log\sigma^{2}-\ell^{2}/\sigma^{2})/2$ to learn $\sigma$. This is equivalent to minimising $\sum\left(\ell^{2}/2/\sigma^{2}+\log\sigma\right)$. When balancing between two losses $\ell_{\text{CLT}}$ and $\ell_{\text{SP}}$ that have deviations $\sigma_{\text{CLT}}$ and $\sigma_{\text{SP}}$, respectively, we need to use Equation (10).
Replacing $\ell_{\text{CLT}}$ in Equation (10) with ${\mathcal{L}}_{\text{CLT}}\left(1-\exp(-{\mathcal{L}}_{\text{CLT}})\right)$ and $\ell_{\text{SP}}$ with ${\mathcal{L}}_{\text{SP}}\left(1-\exp(-{\mathcal{L}}_{\text{SP}})\right)$, Equation (1) is eventually derived to be the complete loss. The training process trades-off between ${\mathcal{L}}_{\text{CLT}}$ and ${\mathcal{L}}_{\text{SP}}$, as well as between the weight regulariser $r_{\bm{\eta}}=\log\sigma_{\text{CLT}}\sigma_{\text{SP}}$ and the rest of Equation (1). When ${\mathcal{L}}_{\text{CLT}}$ is small and ${\mathcal{L}}_{\text{SP}}$ is large, $\sigma_{\text{CLT}}$ becomes small and $\sigma_{\text{SP}}$ becomes large, which then increases the weight for ${\mathcal{L}}_{\text{CLT}}$ while reduces the weight for ${\mathcal{L}}_{\text{SP}}$. The reverse occurs when ${\mathcal{L}}_{\text{CLT}}$ is large and ${\mathcal{L}}_{\text{SP}}$ is small. As the weighted sum of ${\mathcal{L}}_{\text{CLT}}$ and ${\mathcal{L}}_{\text{SP}}$ increases by larger weights, $\log\sigma_{\text{CLT}}\sigma_{\text{SP}}$ decreases and discourages the increase from being too much. Similarly, if the weighted sum decreases by smaller weights, $\log\sigma_{\text{CLT}}\sigma_{\text{SP}}$ also regularises the decrease.